# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Abstract
The emergence of an orbital angular momentum (OAM) response to a charge current holds promise for technological applications, allowing electrical control of magnetization dynamics. Often, the OAM current is invoked in explaining experimental results for very large orbital transport effects, but this is conceptually challenging as the OAM current is not a conserved quantity. Instead of utilizing the orbital conductivity associated with the non-conserved OAM current, we here use non-equilibrium Green’s functions to directly image the OAM density in real space under an applied electric current bias. We find strong spatial variations in OAM density, with the lattice acting as a source and sink of OAM. Moreover, we show that the OAM response depends sensitively on the angle between the charge current and the crystal axis. This enables the generation of an orbital Hall effect in one current direction and solely a longitudinal response in another. We refer to this as an orbital splitter effect, analogous to the spin splitter effect in altermagnets.
@@ I Introduction
Specifically, Ref. [10] demonstrated that an orbital Hall effect (OHE) can emerge from the orbital texture, also in centrosymmetric systems. The analogous effect in spin systems, namely the spin Hall effect (SHE) [11, 12, 13], can also be understood in terms of the spin texture. However, unlike the OHE, the intrinsic SHE relies on SOC and is therefore precluded in centrosymmetric systems. Thus, with the OHE, one is not limited to heavy metals, enabling use of materials that are more beneficial from both a device [14] and an environmental perspective [15]. Furthermore, because the OHE scales with nonrelativistic quantities, it is often estimated to be much larger than its spin counterpart [16, 17, 18]. The OHE was recently observed in Ti [19] and Cr [20] using the magneto-optical Kerr effect (MOKE). As MOKE measures the surface magnetization, these studies offered direct evidence for the OHE, with the caveat that SOC is weak in both materials, making OAM the dominant contributor to the magnetic signal.
An indirect consequence of the OHE is the orbital torque [21], which fulfills the promise of orbitronics regarding the control of magnetization using electric currents. The orbital torque has been observed in a range of different systems [22, 23, 24, 25, 26], but it is generally difficult to disentangle from the spin-orbit torque, especially since orbital-to-spin conversion is believed to significantly boost the orbital torque [27, 21]. This conversion makes multilayer systems an ideal platform for maximizing the orbital torque efficiency, where an orbital current is generated in one layer through the OHE and subsequently converted in an adjacent layer through SOC. The size and efficiency of the orbital torque makes orbital-assisted spin-orbit torque a promising candidate for next-generation memory technology [28].
@@ II System and orbital texture
We will employ a multiorbital tight-binding model with both time-reversal and inversion symmetry. We consider $d$ orbitals belonging to the $t_{2g}$ sector, namely $d_{yz}$, $d_{xz}$, and $d_{xy}$ orbitals, on a square lattice. Associated with each orbital $\alpha$ and lattice site $i$ is the creation and annihilation operators $c_{i\alpha}^{\dagger}$ and $c_{i\alpha}$, such that that a natural basis is $\tilde{c}_{i}=\begin{pmatrix}c_{i,yz}&c_{i,xz}&c_{i,xy}\end{pmatrix}^{\mathrm{T}}$ While these orbitals are, in their own right, $L=2$ orbitals, together they form an effective $L=1$ subspace [38]. Without SOC, both $L$ and spin $S$ remain good and separate quantum numbers. As we focus on OAM and do not consider SOC, the spin degree of freedom can be ignored. Including spin would merely double the size of the basis, without any coupling between the spin up and down block, making computations more costly. Hence, we do not consider the spin degree of freedom, while emphasizing that all our results straightforwardly extend to spinful electrons. Under the assumption of zero SOC, such an extension would simply introduce a spin degeneracy in the energy bands and a doubling of the size of observables.
@@
We work within the atom-centered approximation, only accounting for the OAM associated with the orbitals located on lattice sites [14, 9], neglecting interatomic or itinerant contributions. This is well-justified for systems with inversion symmetry [14] and bulk elemental ferromagnets such as Ni and Co [39]. In both compounds, $t_{2g}$ orbitals play a key role. The associated dimensionless OAM operators $\hat{L}_{l}$, $l=x,y,z$, can be expressed as matrices $\tilde{L}_{l}$
$\tilde{L}_{x}=\begin{pmatrix}0&0&0\\ 0&0&i\\ 0&-i&0\end{pmatrix}\;\tilde{L}_{y}=\begin{pmatrix}0&0&-i\\ 0&0&0\\ i&0&0\end{pmatrix}\;\tilde{L}_{z}=\begin{pmatrix}0&-i&0\\ i&0&0\\ 0&0&0\end{pmatrix},$ (1)
@@
which act on $\tilde{c}_{i}$ and $\tilde{c}^{\dagger}_{i}$. They satisfy the angular momentum algebra $[\tilde{L}_{i},\tilde{L}_{j}]=i\varepsilon_{ijl}\tilde{L}_{l}$. Each $\tilde{L}_{l}$ has the three eigenvalues: $1,0,-1$, corresponding to three eigenvectors that depend on $l$. For concreteness, consider $\tilde{L}_{x}$. A pure $d_{yz}$ state is an eigenvector of $\tilde{L}_{x}$ with eigenvalue 0. The other two eigenvalues of $\tilde{L}_{x}$ are ±$\pm 1$, with corresponding eigenvectors $|\psi_{\pm}\rangle_{x}=d_{xz}\mp id_{xy}$. Thus, it follows that for a state to have nonzero OAM, it must be in the form $(d_{xz}+\mathrm{e}^{i\phi}d_{xy})/\sqrt{2}$, which has $\langle L_{x}\rangle=-\sin\phi$. In other words, a state must be a superposition of orbital states with a relative complex phase factor, breaking time-reversal symmetry, to carry a nonzero OAM. $\tilde{L}_{y}$ and $\tilde{L}_{z}$ have similar structure for the eigenvectors $|\psi_{\pm}\rangle_{y/z}$ and $\langle L_{y/z}\rangle$. Here we used notation from previous works [40, 36].
The Hamiltonian we use was considered in Ref. [40] to study orbital vortices in superconductors. Orbital-preserving hopping processes are consistent with the point group symmetry of the square lattice $C_{4v}$, where we denote associated matrix element as $t_{ij}^{\alpha\beta}$. From symmetry, it then follows that $yz$ and $xz$ hopping may be anisotropic $t_{i,i\pm e_{x}}^{xz,xz}=t_{i,i\pm e_{y}}^{yz,yz}\equiv t$ and $t_{i,i\pm e_{x}}^{yz,yz}=t_{i,i\pm e_{y}}^{xz,xz}\equiv t_{\gamma}$, and that the hopping of $xy$ orbitals is isotropic, $t_{i,i\pm e_{x}}^{xy,xy}=t_{i,i\pm e_{y}}^{xy,xy}\equiv t^{\prime}$. The magnitude of the hopping is determined by the spatial overlap of adjacent atomic orbital wave functions, which is tied to the underlying solid making up the lattice. So while not strictly speaking guaranteed by $C_{4v}$, we set $t^{\prime}=t$ based the spatial overlap of the involved orbitals. Moreover, the spatial overlap of the orbitals causes $t_{\gamma}$ to be smaller than $t$. An orbital hybridization term $t_{m}$ is also included, coupling $d_{yz}$ to $d_{xy}$ in the $y$ direction and $d_{xz}$ to $d_{xy}$ in the $x$ direction: $t_{i,i\pm e_{y}}^{yz,xy}=t_{i,i\pm e_{x}}^{xz,xy}\equiv t_{m}$. There is no coupling directly between $d_{yz}$ and $d_{xz}$ since none of their lobes overlap. An illustration of the hopping processes presented here is given in Fig. 1. In principle, however, $d_{yz}$ and $d_{xz}$ couple to second order in $t_{m}$, mediated by their coupling to $d_{xy}$. Variants of this model, with $t_{m}=0$, has been employed in other works on orbital effects in superconductors [38, 41, 42] and magnetic heterostructures [36].
@@
For $t_{m}=0$, Eq. (3) is already diagonalized; the different $d$ orbitals are eigenvectors corresponding to eigenvalues that can be read of the diagonal of the matrix in Eq. (3). The bandwidth of $\varepsilon_{yz}$ and $\varepsilon_{xz}$ is $4(t+t_{\gamma})$, and for $\varepsilon_{xy}$ it is $8t$. Moreover, in the low-filling limit, it follows immediately that the Fermi surface of $d_{xy}$ is a circle, while for $d_{yz}$ and $d_{xz}$ they are perpendicularly orientated ellipses, with the same eccentricity $\sqrt{1-(t_{\gamma}/t)^{2}}$. The Fermi surface at intermediate filling is illustrated in Fig. 2 (a) and reflects the low-filling behavior.
@@
With hybridization, $t_{m}\neq 0$, $\tilde{H}(\boldsymbol{k})$ must be diagonalized for all $\boldsymbol{k}$. The Fermi surfaces of these bands are shown in Fig. 2 (b), with otherwise equal parameters as in Fig. 2 (a). We label the three eigenvalues of $\tilde{H}(\boldsymbol{k})$ by $\varepsilon_{n}(\boldsymbol{k})$, and in Fig. 2 (b) $\varepsilon_{1}(\boldsymbol{k})$ consists of predominantly $d_{xy}$ orbitals, $\varepsilon_{2}(\boldsymbol{k})$ and $\varepsilon_{3}(\boldsymbol{k})$ are the contours with $d_{yz}$ and $d_{xz}$ character at $k_{y}=0$, respectively. Fig. 2 For $|k_{x}|=|k_{y}|$, we see the avoided crossings between bands $\varepsilon_{2}(\boldsymbol{k})$ and $\varepsilon_{3}(\boldsymbol{k})$ induced by $t_{m}$. As we illustrate, the corresponding eigenvectors $\psi_{1}(\boldsymbol{k})$ and $\psi_{2}(\boldsymbol{k})$ interchange their orbital character, going from predominantly $d_{yz}$ to $d_{xz}$ character and vice versa.
@@
It is first instructive to compare our model to the paradigmatic model introduced in Ref. [10]. There, the two bands contributing to nonzero $L_{z}$ have radial and tangential $p$-orbital character, resulting in nonzero $\sigma_{xy}^{z}$. These bands consist of $p_{x}$ and $p_{y}$ orbitals. The coupling between $p_{x}$ and $p_{y}$ occurs through a second order processes, mediated by an $s$ orbital. Inversion symmetry imposes that hopping between $p_{x/y}$ and $s$ must be odd in $k_{x/y}$, resulting in an effective coupling between $p_{x}$ and $p_{y}$ that is odd in both $k_{x}$ and $k_{y}$. With the $d$ orbitals we employ, the situation is similar for $L_{z}$; two orbitals, $d_{yz}$ and $d_{xz}$ are needed to have nonzero $L_{z}$, and they are only coupled by a second-order process to another orbital, $d_{xy}$. The difference, however, is that the coupling between $d_{x/yz}$ to $d_{xy}$ is even in $k_{x/y}$, rendering the effective coupling between $d_{yz}$ and $d_{xz}$ even in both $k_{x}$ and $k_{y}$.
@@
While the calculation in App. A provides valuable insight into the underlying physics, it has some severe limitations. Specifically, it is only valid for short time scales and can not be used to consider the steady-state of the system. Perhaps more importantly, considering anisotropic hopping $t_{\gamma}\neq t$ precludes a tractable analytical diagonalization of Eq. (3), making it difficult to extract useful information for such systems. These systems are particularly interesting since their anisotropic Fermi surfaces, illustrated in Fig. 2, should manifest in terms of anistropic transport properties. Additionally, because the orbital character of the bands is momentum-dependent, non-equilibrium-induced $L_{l}$ effects should become even more pronounced in the presence of anisotropy. For example, $\varepsilon_{2}(\boldsymbol{k})$ in Fig. 2 (b) exhibits a similar radial character as the band considered in Ref. [10], transitioning from predominantly $d_{yz}$ character near the $x$-axis to $d_{xz}$ character near the $y$-axis. As a last point of note, $t_{\gamma}<t$ is a requirement for studying systems exhibiting similar Fermi surfaces as spin-up and spin-down electrons in effective models of altermagnets, as mentioned earlier.

### v2

@@ Abstract
The emergence of an orbital angular momentum (OAM) response to a charge current holds promise for technological applications, allowing electrical control of magnetization dynamics. Often, the OAM current is invoked in explaining experimental results for very large orbital transport effects, but this is conceptually challenging as the OAM current is not a conserved quantity. Instead of utilizing the orbital conductivity associated with the non-conserved OAM current, we here use non-equilibrium Green’s functions to directly image the OAM density in real space under an applied electric current bias. We find strong spatial variations in OAM density, with the lattice acting as a source and sink of OAM. Moreover, we show that the OAM response depends sensitively on the angle between the charge current and the crystal axis. This enables the generation of a transverse OAM response in one current direction and solely a longitudinal response in another. We refer to this as an orbital splitter effect, analogous to the spin splitter effect in altermagnets.
@@ I Introduction
Specifically, Ref. [10] demonstrated that an orbital Hall effect (OHE) can emerge from the orbital texture, also in centrosymmetric systems. The analogous effect in spin systems, namely the spin Hall effect (SHE) [11, 12, 13], can also be understood in terms of the spin texture. However, unlike the OHE, the intrinsic SHE relies on SOC and is therefore precluded in centrosymmetric systems. Thus, with the OHE, one is not limited to heavy metals, enabling use of materials that are more beneficial from both a device [14] and an environmental perspective [15]. Furthermore, because the OHE scales with nonrelativistic quantities, it is often estimated to be much larger than its spin counterpart [16, 17, 18]. The OHE was recently observed in Ti [19] and Cr [20] using the magneto-optical Kerr effect (MOKE). The orbital Hall effect and orbital magnetization induced by an electric field have also been studied [21, 22, 23, 24, 25] in transition metal dichalcogenides. As MOKE measures the surface magnetization, these studies offered direct evidence for the OHE, with the caveat that SOC is weak in both materials, making OAM the dominant contributor to the magnetic signal.
An indirect consequence of the OHE is the orbital torque [26], which fulfills the promise of orbitronics regarding the control of magnetization using electric currents. The orbital torque has been observed in a range of different systems [27, 28, 29, 30, 31], but it is generally difficult to disentangle from the spin-orbit torque, especially since orbital-to-spin conversion is believed to significantly boost the orbital torque [32, 26]. This conversion makes multilayer systems an ideal platform for maximizing the orbital torque efficiency, where an orbital current is generated in one layer through the OHE and subsequently converted in an adjacent layer through SOC. The size and efficiency of the orbital torque make orbital-assisted spin-orbit torque a promising candidate for next-generation memory technology [33].
@@ II System and orbital texture
We employ a multiorbital tight-binding model and consider $d$ orbitals belonging to the $t_{2g}$ sector, namely $d_{yz}$, $d_{xz}$, and $d_{xy}$ orbitals, on a square lattice. Associated with each orbital $\alpha$ and lattice site $i$ are the creation and annihilation operators $c_{i\alpha}^{\dagger}$ and $c_{i\alpha}$, such that that a natural basis is $\tilde{c}_{i}=\begin{pmatrix}c_{i,yz}&c_{i,xz}&c_{i,xy}\end{pmatrix}^{\mathrm{T}}$ While these orbitals are, in their own right, $L=2$ orbitals, together they form an effective $L=1$ subspace [43]. Without SOC, both $L$ and spin $S$ remain good and separate quantum numbers. As we focus on OAM and do not consider SOC, the spin degree of freedom can be ignored. Including spin would merely double the size of the basis, without any coupling between the spin up and down block, making computations more costly. Hence, we do not consider the spin degree of freedom, while emphasizing that all our results straightforwardly extend to spinful electrons. Under the assumption of zero SOC, such an extension would simply introduce a spin degeneracy in the energy bands and a doubling of the size of observables.
@@
We work within the atom-centered approximation, only accounting for the OAM associated with the orbitals located on lattice sites [14, 9], neglecting interatomic or itinerant contributions. This is well-justified, for example, in elemental ferromagnets such as Ni and Co [44]. In both compounds, $t_{2g}$ orbitals play a key role. The associated dimensionless OAM operators $\hat{L}_{l}$, $l=x,y,z$, can be expressed as matrices $\tilde{L}_{l}$
$\tilde{L}_{x}=\begin{pmatrix}0&0&0\\ 0&0&i\\ 0&-i&0\end{pmatrix}\;\tilde{L}_{y}=\begin{pmatrix}0&0&-i\\ 0&0&0\\ i&0&0\end{pmatrix}\;\tilde{L}_{z}=\begin{pmatrix}0&i&0\\ -i&0&0\\ 0&0&0\end{pmatrix},$ (1)
@@
which act on $\tilde{c}_{i}$ and $\tilde{c}^{\dagger}_{i}$. We will use the tilde notation whenever the specific matrix representation of the OAM operators in Eq. (1) is concerned, using $\hat{L}_{l}$ or $L_{l}$ otherwise. $\tilde{L}_{l}$ satisfy a modified angular momentum algebra, $[\tilde{L}_{i},\tilde{L}_{j}]=-i\varepsilon_{ijl}\tilde{L}_{l}$, where the additional minus sign arises from truncating the full $d$-orbital space [45]. Each $\tilde{L}_{l}$ has the three eigenvalues: $1,0,-1$, corresponding to three eigenvectors that depend on $l$. For concreteness, consider $\tilde{L}_{x}$. A pure $d_{yz}$ state is an eigenvector of $\tilde{L}_{x}$ with eigenvalue 0. The other two eigenvalues of $\tilde{L}_{x}$ are ±$\pm 1$, with corresponding eigenvectors $|\psi_{\pm}\rangle_{x}=d_{xz}\mp id_{xy}$. Thus, it follows that for a state to have nonzero OAM, it must be in the form $(d_{xz}+\mathrm{e}^{i\phi}d_{xy})/\sqrt{2}$, which has $\langle L_{x}\rangle=-\sin\phi$. In other words, a state must be a superposition of orbital states with a relative complex phase factor, breaking time-reversal symmetry, to carry a nonzero OAM. $\tilde{L}_{y}$ and $\tilde{L}_{z}$ have similar structure for the eigenvectors $|\psi_{\pm}\rangle_{y/z}$ and $\langle L_{y/z}\rangle$. Here we used notation from previous works [46, 41].
The Hamiltonian we use was considered in Ref. [46] to study orbital vortices in superconductors. Orbital-preserving hopping processes are made consistent with the point group symmetry of the square lattice $C_{4v}$, where we denote the associated matrix element as $t_{ij}^{\alpha\beta}$. Similar to Refs. [46, 47], we also include an orbital hybridization term $t_{m}$, coupling $d_{yz}$ to $d_{xy}$ in the $y$ direction and $d_{xz}$ to $d_{xy}$ in the $x$ direction: $t_{i,i\pm e_{y}}^{yz,xy}=t_{i,i\pm e_{x}}^{xz,xy}\equiv t_{m}$, which breaks multiple symmetries [46]. They break all mirror symmetries and thus also inversion symmetry, as well as fourfold symmetry. Out-of-plane symmetry can be broken by interfacing the two-dimensional layer with a substrate and/or vacuum. The breaking of the remaining symmetries is more complicated as $t_{m}$ is not consistent with the point group symmetry of the square lattice. The lattice must thus be distorted for $t_{m}$ to be nonzero.
However, we still keep to modeling the system as an effective square lattice. Importantly, we do so under the assumption that $t_{m}$ arises from an underlying crystal structure, with lower point-group symmetry than the square lattice. A nonzero $t_{m}$ is then permissible due to, for example, bonding between the square lattice sites and ligand atoms or by introducing inhomogeneous strain [48], breaking the aforementioned symmetries. The breaking of these symmetries also affects the $d$ orbitals directly, but as we work within the tight-binding approximation, we will not consider this in detail. Instead, we let $t_{m}$ capture the effects from the underlying symmetry-breaking crystal field. We note that including a term ($t_{m}$) which breaks symmetries not present in the lattice itself (square) is a standard treatment in the literature, for instance in square lattices with Rashba spin-orbit interactions. Two-dimensional noncentrosymmetric oxide heterostructures grown along the [111] crystallographic orientation have been suggested as candidate systems with the appropriate crystal symmetry to host $t_{m}\neq 0$ [47, 48].
An illustration of the hopping processes presented here is given in Fig. 1. As several symmetries must be broken for $t_{m}\neq 0$, it is worth dwelling on the relative sizes of hopping magnitudes and other hopping processes allowed by symmetry. The orbital-preserving processes are kept consistent with the square lattice point group symmetry, as $d_{yz}$ and $d_{xz}$ hoppings are anisotropic $t_{i,i\pm e_{x}}^{xz,xz}=t_{i,i\pm e_{y}}^{yz,yz}\equiv t$ and $t_{i,i\pm e_{x}}^{yz,yz}=t_{i,i\pm e_{y}}^{xz,xz}\equiv t_{\gamma}$ and the hopping of $d_{xy}$ orbitals is isotropic $t_{i,i\pm e_{x}}^{xy,xy}=t_{i,i\pm e_{y}}^{xy,xy}\equiv t^{\prime}$. Because the magnitude of the hopping is determined by the spatial overlap of adjacent atomic orbital wave functions, we take $t^{\prime}=t$ (which we use as the base energy unit throughout the paper) and consider $t_{\gamma}<t$. The difference in spatial overlaps is also the reason why the $t_{m}$ process is included, while there is no direct coupling between $d_{yz}$ and $d_{xz}$: the overlap between $d_{xy}$ and the $d_{yz}$ and $d_{xz}$ orbitals is simply larger than that of $d_{yz}$ and $d_{xz}$ since the latter do not have overlapping lobes. Within our model, $d_{yz}$ and $d_{xz}$ still couple to second order in $t_{m}$, mediated by their coupling to $d_{xy}$.
Finally, with the breaking of symmetries required for $t_{m}\neq 0$, the orbital Rashba coupling $t_{\text{SO}}$ can also be nonzero. In orbitronics, $t_{\text{SO}}$ is known to generate both orbital Hall and Edelstein responses [49]. In our particular model, the coupling takes the form $t_{\text{SO}}(\sin k_{y}\tilde{L}_{x}-\sin k_{x}\tilde{L}_{y})$ in momentum space [50]. While symmetry thus allows $t_{\text{SO}}\neq 0$ whenever $t_{m}\neq 0$, because its associated effects are well-studied, we will mainly focus on the physics induced by $t_{m}\neq 0$, putting $t_{\text{SO}}$ to zero in most of our results. Still, to connect our framework to previous findings, we will also include some results with $t_{\text{SO}}\neq 0$ when explicitly stating so.
@@
For $t_{m}=0$, Eq. (3) is already diagonalized; the different $d$ orbitals are eigenvectors corresponding to eigenvalues that can be read off from the diagonal of the matrix in Eq. (3). The bandwidth of $\varepsilon_{yz}$ and $\varepsilon_{xz}$ is $4(t+t_{\gamma})$, and for $\varepsilon_{xy}$ it is $8t$. Moreover, in the low-filling limit, it follows immediately that the Fermi surface of $d_{xy}$ is a circle, while for $d_{yz}$ and $d_{xz}$ they are perpendicularly orientated ellipses, with the same eccentricity $\sqrt{1-(t_{\gamma}/t)^{2}}$. The Fermi surface at intermediate filling is illustrated in Fig. 2 (a) and reflects the low-filling behavior.
@@
With hybridization, $t_{m}\neq 0$, $\tilde{H}(\boldsymbol{k})$ must be diagonalized for all $\boldsymbol{k}$. The Fermi surfaces of these bands are shown in Fig. 2 (b), with otherwise equal parameters as in Fig. 2 (a). We label the three eigenvalues of $\tilde{H}(\boldsymbol{k})$ by $\varepsilon_{n}(\boldsymbol{k})$, and in Fig. 2 (b) $\varepsilon_{1}(\boldsymbol{k})$ consists of predominantly $d_{xy}$ orbitals, $\varepsilon_{2}(\boldsymbol{k})$ and $\varepsilon_{3}(\boldsymbol{k})$ are the contours with $d_{yz}$ and $d_{xz}$ character at $k_{y}=0$, respectively. For $|k_{x}|=|k_{y}|$, we see the avoided crossings between bands $\varepsilon_{2}(\boldsymbol{k})$ and $\varepsilon_{3}(\boldsymbol{k})$ induced by $t_{m}$. As we illustrate, the corresponding eigenvectors $\psi_{1}(\boldsymbol{k})$ and $\psi_{2}(\boldsymbol{k})$ interchange their orbital character, going from predominantly $d_{yz}$ to $d_{xz}$ character and vice versa. Both for the sake of completeness and later relevance, the band structure for a system with higher $t_{m}=0.2$ is given in Fig. 3.
Figure 3: The energy bands for $H(\boldsymbol{k})$ in Eq. (3) with $t_{m}=0.2$ and $t_{\gamma}=0.5$ along high-symmetry points. Inset: Showing the small $\boldsymbol{k}$ behavior near $\Gamma$ with dashed lines showing the two values for $\mu$ used in Fig. 6.
@@
It is first instructive to compare our model to the paradigmatic model introduced in Ref. [10]. There, the two bands contributing to nonzero $L_{z}$ have radial and tangential $p$-orbital character, resulting in nonzero $\sigma_{xy}^{z}$. These bands consist of $p_{x}$ and $p_{y}$ orbitals. The coupling between $p_{x}$ and $p_{y}$ occurs through a second order processes, mediated by an $s$ orbital. Because Ref. [10] has inversion symmetry, it imposes that hopping between $p_{x/y}$ and $s$ must be odd in $k_{x/y}$, resulting in an effective coupling between $p_{x}$ and $p_{y}$ that is odd in both $k_{x}$ and $k_{y}$. With the $d$ orbitals we employ, the situation is similar for $L_{z}$, albeit without inversion symmetry; two orbitals, $d_{yz}$ and $d_{xz}$ are needed to have nonzero $L_{z}$, and they are only coupled by a second-order process to another orbital, $d_{xy}$. The difference, however, is that the coupling between $d_{x/yz}$ to $d_{xy}$ is even in $k_{x/y}$, rendering the effective coupling between $d_{yz}$ and $d_{xz}$ even in both $k_{x}$ and $k_{y}$. As we will now see, this difference in the coupling of orbitals leads to a different OAM response to a charge current than the OHE observed in the $sp$-model of Ref. [10].
@@
While the calculation in App. A provides valuable insight into the underlying physics, it has some severe limitations. Specifically, it is only valid for short time scales and can not be used to consider the steady-state of the system. Perhaps more importantly, considering anisotropic hopping $t_{\gamma}\neq t$ precludes a tractable analytical diagonalization of Eq. (3), making it difficult to extract useful information for such systems. These systems are particularly interesting since their anisotropic Fermi surfaces, illustrated in Fig. 2, should manifest in terms of anistropic transport properties. Additionally, because the orbital character of the bands is momentum-dependent, non-equilibrium-induced OAM effects should become even more pronounced in the presence of anisotropy. For example, $\varepsilon_{2}(\boldsymbol{k})$ in Fig. 2 (b) exhibits a similar radial character as the band considered in Ref. [10], transitioning from predominantly $d_{yz}$ character near the $x$-axis to $d_{xz}$ character near the $y$-axis. As a last point of note, $t_{\gamma}<t$ is a requirement for studying systems exhibiting similar Fermi surfaces as spin-up and spin-down electrons in effective models of altermagnets, as mentioned earlier.